OpenStudy (anonymous):

A braking car on a drying pavement has a constant acceleration of 3.0 m/s^2. If the car was moving at 10.0 m/s, approximately how long would it take to stop? 10 s, 17 s, 30 s, or 3.3 s @jim_thompson5910 :)

OpenStudy (anonymous):

so a=3 t=? vi=0 and would vf=10 ? @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

"If the car was moving at 10.0 m/s, approximately how long would it take to stop?" the car is initially going 10 m/s (vi) the stopped speed is 0 m/s (vf)

OpenStudy (anonymous):

ohh okay so vi=10 vf=0 a=3 t=? so using vf=vi+at again?

OpenStudy (anonymous):

0=10+3t ?

jimthompson5910 (jim_thompson5910):

yes, solve for t

OpenStudy (anonymous):

3t=-10 t=-3.3333.... ?

jimthompson5910 (jim_thompson5910):

I forgot to mention, but since you're decelerating, the value of 'a' is negative

jimthompson5910 (jim_thompson5910):

a = -3

OpenStudy (anonymous):

ohhh okay:) so 0=10+(-3t) -10=-3t t=3.3333 so 3.3 s=t ?

jimthompson5910 (jim_thompson5910):

yes roughly

OpenStudy (anonymous):

and so anytime it's decelerating, it will be -? and okie yay!

jimthompson5910 (jim_thompson5910):

yes 'a' is negative when you decelerate

OpenStudy (anonymous):

awesome!!! thanks bunches!! :)